# Theoretical Insights into Overparameterized Models in Multi-Task and Replay-Based Continual Learning

### v1

@@ C.2 Proof of Theorems
where $\hat{X}^{\dagger}_{1t}=\hat{X}_{1t}(\hat{X}^{T}_{1t}\hat{X}_{1t})^{-1}$, $\hat{P}_{1t}=\hat{X}_{1t}(\hat{X}^{T}_{1t}\hat{X}_{1t})^{-1}\hat{X}^{T}_{1t}$ and $\hat{P}_{0s}=\hat{X}_{1t}(\hat{X}^{T}_{1t}\hat{X}_{1t})^{-1}\hat{X}^{T}_{0s}$. Notice that the first $tm$ columns of $\hat{X}_{1t}$ is exactly $X_{1t}$ and there are exactly $n-m$ extra columns in $\hat{X}_{1t}$ that we threw away before proceeding to the task $t+1$. Let $\hat{X}_{t}\in\mathbb{R}^{p\times(n-m)}$ represent this portion of the training set. In other words, $\hat{X}_{1t}=\begin{bmatrix}X_{1t}&\hat{X}_{t}\end{bmatrix}$. With this in mind, based on Lemma C.3 part (v), it holds that
$\hat{X}^{\dagger}_{1t}=\begin{bmatrix}X^{\dagger}_{1t}-(I-P_{1t})\hat{X}_{t}(\hat{X}_{t}^{T}(I-P_{1t})\hat{X}_{t})^{-1}\hat{X}_{t}^{T}X^{\dagger}_{1t}&(I-P_{1t})\hat{X}_{t}(\hat{X}_{t}^{T}(I-P_{1t})\hat{X}_{t})^{-1}\end{bmatrix}$
$\Rightarrow P_{1t}\hat{P}_{1t}=P_{1t}\hat{X}^{\dagger}_{1t}\hat{X}^{T}_{1t}=\begin{bmatrix}X^{\dagger}_{1t}&0\end{bmatrix}\hat{X}^{T}_{1t}=P_{1t}$
@@
$\mathbb{E}[\|P_{1t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$ $=\mathbb{E}[\|P_{1t}(I-\hat{P}_{1t})(w_{{\bm{v}}{t-1}}-w^{*}_{i})+\sum_{s=1}^{t}P_{1t}\hat{P}_{0s}(w_{s}^{*}-w^{*}_{i})+P_{1t}\hat{X}^{\dagger}_{1t}\hat{z}_{1t}\|^{2}]$
$=\mathbb{E}[\|\sum_{s=1}^{t}P_{0s}(w_{s}^{*}-w^{*}_{i})+X^{\dagger}_{1t}z_{1t}\|^{2}]$
We have previously calculated terms like this in Section C.2.2. Therefore, we can write overall:
$\mathbb{E}[\|(I-P_{1r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$ $=\mathbb{E}[\|w_{{\bm{v}}{t}}-w^{*}_{i}\|^{2}]-(1-\frac{n}{p-tm})\mathbb{E}[\|P_{1t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]-(\frac{n}{p-tm})\mathbb{E}[\|w_{{\bm{v}}{t}}-w^{*}_{i}\|^{2}]$
$=\alpha_{t}\mathbb{E}[\|w_{{\bm{v}}{t}}-w^{*}_{i}\|^{2}]-\alpha_{t}\mathbb{E}[\|P_{1t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$
@@
$+\frac{m}{p}\sum_{s=1}^{t}\|w^{*}_{s}-w_{i}^{*}\|^{2}+\frac{n}{p}\|w^{*}_{t+1}-w_{i}^{*}\|^{2}+\mathbb{E}[\|(I-P_{1r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+\frac{\bar{n}_{t}\sigma^{2}}{p-\bar{n}_{t}-1}$
@@
Using the recursive form in Equation 24, one can exactly find the loss of a replay+regularization-based continual learner. However, the full form is not intuitive. Therefore, we focus on the two-task case instead.
@@
Appendix D Effect of Regularization in CL Models
In this section, we study the effect of regularization on the performance of replay-based continual models by using the results of Theorems 4.1 and 4.2. To that end, let’s reconsider the Equations 14 to 17 and denote the $G$ terms in these equations as $G^{\text{rep}}$ and $G^{\text{reg}}$ to respectively represent the generalization error of the pure replay-based and regularization+replay CL models. We apply similar notations to the $F$ terms to have $F^{\text{rep}}$ and $F^{\text{reg}}$ which lets us write
$G^{\text{rep}}=\underbrace{\frac{n}{2p}(1+\frac{m}{n}+\frac{2m}{p-(n+m)-1})\|w_{1}^{*}-w_{2}^{*}\|_{2}^{2}}_{\text{term $G^{\text{rep}}_{1}$}}+\underbrace{(1-\frac{n+m}{p})\|w_{1}^{*}\|_{2}^{2}}_{\text{term $G^{\text{rep}}_{2}$}},$
$F^{\text{rep}}=\underbrace{\frac{n}{p}(1+\frac{m}{p-(n+m)-1)}\|w_{1}^{*}-w_{2}^{*}\|_{2}^{2}}_{\text{term $F^{\text{rep}}_{1}$}}-\underbrace{\frac{m}{p}\|w_{1}^{*}\|_{2}^{2}}_{\text{term $F^{\text{rep}}_{2}$}},$
$G^{\text{reg}}=\underbrace{\frac{n}{2p}(2-\frac{n-m}{p-m}+\frac{2m}{p-(n+m)-1})\|w_{1}^{*}-w_{2}^{*}\|_{2}^{2}}_{\text{term $G^{\text{reg}}_{1}$}}+\underbrace{(1-\frac{n}{p-m})(1-\frac{n}{p})\|w_{1}^{*}\|_{2}^{2}}_{\text{term$G^{\text{reg}}_{2}$}},$
$F^{\text{reg}}=\underbrace{\frac{n}{p}(1+\frac{m}{p-(n+m)-1})\|w_{1}^{*}-w_{2}^{*}\|_{2}^{2}}_{\text{term $F^{\text{reg}}_{1}$}}-\underbrace{\frac{n}{p-m}(1-\frac{n}{p})\|w_{1}^{*}\|_{2}^{2}}_{\text{term $F^{\text{reg}}_{2}$}}.$
We start by comparing the forgetting terms:
$F^{\text{rep}}-F^{\text{reg}}$ $=F^{\text{rep}}_{1}+F^{\text{rep}}_{2}-F^{\text{reg}}_{1}-F^{\text{reg}}_{2}$
$=-\frac{m}{p}\|w_{1}^{*}\|_{2}^{2}+\frac{n}{p-m}(1-\frac{n}{p})\|w_{1}^{*}\|_{2}^{2}$
$=\frac{1}{p(p-m)}(-m(p-m)+n(p-n))\|w_{1}^{*}\|_{2}^{2}$
$=\frac{(n-m)}{p(p-m)}(p-(n+m))\|w_{1}^{*}\|_{2}^{2}>0$
Noticing that $p>n+m$ and $m\leq n$, reveals that always $F^{\text{rep}}>F^{\text{reg}}$ which means that the regularized continual learner always has lower forgetting. Similarly, we can compare the $G$ terms:
$G^{\text{rep}}_{1}-G^{\text{reg}}_{1}$ $=\frac{n}{2p}(-1+\frac{m}{n}+\frac{n-m}{p-m})\|w_{1}^{*}-w_{2}^{*}\|_{2}^{2}$
$=\frac{-(n-m)}{2p(p-m)}(p-m-n)\|w_{1}^{*}-w_{2}^{*}\|_{2}^{2}<0$
This shows that the first generalization error is better for the unregularized continual learner. For the second term,
$G^{\text{rep}}_{2}-G^{\text{reg}}_{2}$ $=(1-\frac{n+m}{p}-(1-\frac{n}{p-m})(1-\frac{n}{p}))\|w_{1}^{*}\|_{2}^{2}$
$=(-\frac{m}{p}+\frac{n}{p-m}-\frac{n}{p-m}\frac{n}{p})\|w_{1}^{*}\|_{2}^{2}$
$=\frac{1}{p(p-m)}(-m(p-m)+np-n^{2})\|w_{1}^{*}\|_{2}^{2}$
$=\frac{n-m}{p(p-m)}(p-(m+n))\|w_{1}^{*}\|_{2}^{2}>0$
This difference is a positive value since $p>m+n$ and $n\geq m$. Therefore, for the second term, the regularization does not benefit. With this in mind, we can not clearly express whether the regularization is helpful in the overall generalization error of the continual learner and it depends on the task similarity value $\|w_{1}^{*}-w_{2}^{*}\|_{2}^{2}$ to decide on the effectiveness of the regularization. Figure 7 provides a visual comparison of the regularized and unregularized memory replay continual learners.
(a) collaborative tasks
(b) conflicting tasks
(c) collaborative tasks
(d) conflicting tasks
Figure 7: Plotting average generalization error of the replay-based and replay+regularization continual learning methods w.r.t. model parameters $p$. For all plots, $T=10$, $\sigma=0$, and for all tasks $n_{t}=50$ and $\|w_{t}^{*}\|^{2}=1$. (a) and (b) show the generalization error ($G$) of a replay-based continual learner for different values of the buffer size ($m$). (c) and (d) are similar plots for the replay+regularization continual learning method. $m=0$ corresponds to using no memory and $m=50$ means using a multi-task learner with maximum memory capacity. In subfigures (a) and (c), the tasks are designed to be collaborative by adjusting $\langle w^{*}_{i},w^{*}_{j}\rangle=\cos\frac{\pi}{8}$ for every pair of task vectors $w^{*}_{i}$ and $w^{*}_{j}$. For subfigures (b) and (d), the tasks are chosen to be conflicting by considering $\langle w^{*}_{i},w^{*}_{j}\rangle=\cos\frac{7\pi}{8}$. Solid line curves represent the results from theoretical predictions. The dot marks are the empirical evaluations averaged over 500 repetitions and are perfectly aligned with the theoretical results in the overparameterized region.
Appendix E Bridging Linear Models and Overparameterized DNNs
@@
$f(z;\theta)\approx f(z;\theta_{0})+\nabla_{\theta}f(z;\theta_{0})^{T}(\theta-\theta_{0})\approx\nabla_{\theta}f(z;\theta_{0})^{T}(\theta-\theta_{0}).$
@@
Additionally, this approximation allows us to define a Neural Tangent Kernel (NTK) (Jacot et al., 2018) as $\Theta(z_{i},z_{j})=\nabla_{\theta}f(z_{i};\theta_{0})^{T}\nabla_{\theta}f(z_{j};\theta_{0})$, which captures the inner product of the gradients of the neural network with respect to the parameters at different data points. As training progresses, if the parameters remain close to their initialization (lazy training), the predictions of the neural network can be effectively described by a linear model in this high-dimensional feature space defined by the NTK. This connection elucidates how overparameterized deep neural networks can exhibit behavior akin to kernel methods, where the NTK serves as the kernel function that determines the similarity between data points.
Appendix F Experimental Details
@@
Appendix G Additional Experiments
In this section, we provide extra experiments on DNNs. Section G.1 provides the MTL experiments with single-head and multi-head architectures but with a frozen backbone. Section G.2 offers the results for the experiments with unfrozen backbones. In Section G.3, we dig deeper into investigating the optimal weights learned by each MLP layer, given that the layers are unshared across the tasks, and we measure the task similarity at each layer using the average inner product of the optimal task-specific weights. In Section G.4, we investigate the effect of the MLP depth and provide the corresponding train and test curves. Finally, in Section G.5 we offer extra experiments on CL models and demonstrate the effect of label noise on CL models.
G.1 Complementary Multi-Task Learning Experiments
Figures 8 to 14 demonstrate the full training and testing curves for training the model in an MTL setup with multi-head and single-head architectures with different backbones and datasets. As discussed before in the main text, the training error is zero for both cases in large enough models. However, the multi-task learner performs worse in the single-head architecture due to the conflict at the classification layer.
Figure 8: Comparing the single-task and multi-task performance on CIFAR-100 with pretrained ResNet-18 backbone. The top row shows the experiment with the single-head architecture, and the bottom row corresponds to the multi-head architecture. The horizontal axis in all plots is $\log_{2}(k)$ where $k$ represents the width of the MLP layer which is placed on top of the feature extractor. $\sigma$ represents the noisy portion of the training set that was corrupted by randomly switching its labels.
@@ G.1 Complementary Multi-Task Learning Experiments
G.2 Multi-Task Learning Experiments with Unfrozen Backbone
@@
Figure 15: Comparing the single-task and multi-task performance on CIFAR-100 with unfrozen ResNet-18 backbone with random initialization. The top row shows the experiment with the single-head architecture, and the bottom row corresponds to the multi-head architecture. The horizontal axis in all plots is $\log_{10}(k)$ where $k$ controls the number of filters in convolutional layers. $\sigma$ represents the noisy portion of the training set that was corrupted by randomly switching its labels.
G.3 Investigation of the Task Optimal Weights
@@
Figure 16: Investigating the effect of weight sharing in MTL by unsharing different layers of the MLP layer and training task-specific layers. The horizontal axis is $\log_{2}(k)$ where $k$ represents the width of the MLP layer.
Figure 17: Comparing the average inner product score (Equation 25) across different layers after unsharing the weights across the tasks. The horizontal axis is $\log_{2}(k)$ where $k$ represents the width of the MLP layer.
G.4 Investigation of the Depth
In this section, we investigate the effect of the MLP depth on top of the pretrained backbones. Figure 18 compares different models with varying depths of the MLP. Notice that our theories are based on the assumption of a linear model, and therefore, they do not directly predict the behavior as a function of the depth. However, as discussed in Section E, the DNN performs similarly to a linear model as long as certain conditions hold. Figure 18 confirms this claim by showing similar behavior for all different models. Another interesting observation in this plot is the impact of depth over the height of the test error peak, which is a notable direction for future studies.
Figure 18: Visualizing the effect of the depth on the performance curves of the multi-head multi-task learner. The horizontal axis is $\log_{2}(k)$ where $k$ represents the width of the MLP layers. Notice that a simple linear layer on top of the feature extractor can not fully overfit the training samples and has no tunable hyperparameter for increasing the complexity.
G.5 Complementary Continual Learning Experiments
@@
Figure 19: Visualizing the CL performance on CIFAR-100 with different values of memory budget denoted as $m$. The horizontal axis is $\log_{2}(k)$ where $k$ represents the width of the MLP layer. The top row corresponds to the experiments with a single-head classifier and the bottom row shows the multi-head architecture.
Figure 20: Visualizing the CL performance on CIFAR-100 using an unfrozen ResNet-18 backbone and a multi-head classification head. The plot shows the performance for different values of memory budget denoted as $m$. The horizontal axis is $\log_{10}(k)$ where $k$ controls the number of filters in the ResNet-18 backbone.
@@
Figure 21: Visualizing the CL performance of multi-head models using memory size $m=50$ on CIFAR-100 with different levels of noise denoted as $\sigma$. The horizontal axis is $\log_{2}(k)$ where $k$ represents the width of the MLP layer.
Appendix H Limitations and Future Work
@@
The next potential limitation is that our analysis is focused on linear models. Although we connected the linear models and DNNs in Appendix E, but the connection holds under certain assumptions which may be violated in different applications. Besides, we emphasize that the goal of this work is not to make formal claims regarding the performance of multi-task DNNs, rather, we tried to qualitatively connect our understanding of linear models to the observations derived experimentally from DNNs. We believe that knowing the trade-offs between different system specifications, such as the model size and number of training samples, can provide an important guideline when deploying practical models.
@@
Appendix I Infrastructures, Code and Data Availability
Related source code is submitted accordingly. In this repository, there exists a README file containing instructions and configuration details. We mainly used PyTorch (Paszke et al., 2019) for our implementations and datasets are freely accessible (Imagenet-R (Hendrycks et al., 2021), CUB(Wah et al., 2011), CIFAR-100 (Krizhevsky et al., 2009), and MNIST (Deng, 2012)). All experiments in the main text are reproducible on a single NVIDIA 2080 TI GeForce GPU.

### v2

@@ A.2 Proof of Theorems
where $\hat{X}^{\dagger}_{1:t}=\hat{X}_{1:t}(\hat{X}^{\top}_{1:t}\hat{X}_{1:t})^{-1}$, $\hat{P}_{1:t}=\hat{X}_{1:t}(\hat{X}^{\top}_{1:t}\hat{X}_{1:t})^{-1}\hat{X}^{\top}_{1:t}$ and $\hat{P}_{0|s}=\hat{X}_{1:t}(\hat{X}^{\top}_{1:t}\hat{X}_{1:t})^{-1}\hat{X}^{\top}_{0|s}$. Notice that the first $tm$ columns of $\hat{X}_{1:t}$ is exactly $X_{1:t}$ and there are exactly $n-m$ extra columns in $\hat{X}_{1:t}$ that we threw away before proceeding to the task $t+1$. Let $\hat{X}_{t}\in\mathbb{R}^{p\times(n-m)}$ represent this portion of the training set. In other words, $\hat{X}_{1:t}=\begin{bmatrix}X_{1:t}&\hat{X}_{t}\end{bmatrix}$. With this in mind, based on Lemma A.3 part (v), it holds that
$\hat{X}^{\dagger}_{1:t}=\begin{bmatrix}X^{\dagger}_{1:t}-(I-P_{1:t})\hat{X}_{t}(\hat{X}_{t}^{\top}(I-P_{1:t})\hat{X}_{t})^{-1}\hat{X}_{t}^{\top}X^{\dagger}_{1:t}&(I-P_{1:t})\hat{X}_{t}(\hat{X}_{t}^{\top}(I-P_{1:t})\hat{X}_{t})^{-1}\end{bmatrix}$
$\Rightarrow P_{1:t}\hat{P}_{1:t}=P_{1:t}\hat{X}^{\dagger}_{1:t}\hat{X}^{\top}_{1:t}=\begin{bmatrix}X^{\dagger}_{1:t}&0\end{bmatrix}\hat{X}^{\top}_{1:t}=P_{1:t}$
@@
$\mathbb{E}[\|P_{1:t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$ $=\mathbb{E}[\|P_{1:t}(I-\hat{P}_{1:t})(w_{{\bm{v}}{t-1}}-w^{*}_{i})+\sum_{s=1}^{t}P_{1:t}\hat{P}_{0|s}(w_{s}^{*}-w^{*}_{i})+P_{1:t}\hat{X}^{\dagger}_{1:t}\hat{z}_{1:t}\|^{2}]$
$=\mathbb{E}[\|\sum_{s=1}^{t}P_{0|s}(w_{s}^{*}-w^{*}_{i})+X^{\dagger}_{1:t}z_{1:t}\|^{2}]$
We have previously calculated terms like this in Section A.2.2. Therefore, we can write overall:
$\mathbb{E}[\|(I-P_{1:r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$ $=\mathbb{E}[\|w_{{\bm{v}}{t}}-w^{*}_{i}\|^{2}]-(1-\frac{n}{p-tm})\mathbb{E}[\|P_{1:t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]-(\frac{n}{p-tm})\mathbb{E}[\|w_{{\bm{v}}{t}}-w^{*}_{i}\|^{2}]$
$=\alpha_{t}\mathbb{E}[\|w_{{\bm{v}}{t}}-w^{*}_{i}\|^{2}]-\alpha_{t}\mathbb{E}[\|P_{1:t}(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]$
@@
$+\frac{m}{p}\sum_{s=1}^{t}\|w^{*}_{s}-w_{i}^{*}\|^{2}+\frac{n}{p}\|w^{*}_{t+1}-w_{i}^{*}\|^{2}+\mathbb{E}[\|(I-P_{1:r})(w_{{\bm{v}}{t}}-w^{*}_{i})\|^{2}]+\frac{\bar{n}_{t}\sigma^{2}}{p-\bar{n}_{t}-1}$
@@
Using the recursive form in Equation 23, one can exactly find the loss of the continual learner in Equation 16. However, the full form is not intuitive. Therefore, we focus on the two-task case instead.
@@
Appendix B Bridging Linear Models and Overparameterized DNNs
@@
$f(z;\theta)\approx f(z;\theta_{0})+\nabla_{\theta}f(z;\theta_{0})^{\top}(\theta-\theta_{0})\approx\nabla_{\theta}f(z;\theta_{0})^{\top}(\theta-\theta_{0}).$
@@
Additionally, this approximation allows us to define a Neural Tangent Kernel (NTK) (Jacot et al., 2018) as $\Theta(z_{i},z_{j})=\nabla_{\theta}f(z_{i};\theta_{0})^{\top}\nabla_{\theta}f(z_{j};\theta_{0})$, which captures the inner product of the gradients of the neural network with respect to the parameters at different data points. As training progresses, if the parameters remain close to their initialization (lazy training), the predictions of the neural network can be effectively described by a linear model in this high-dimensional feature space defined by the NTK. This connection elucidates how overparameterized deep neural networks can exhibit behavior akin to kernel methods, where the NTK serves as the kernel function that determines the similarity between data points.
Appendix C Experimental Details
@@
Appendix D Additional Experiments
In this section, we provide extra experiments on DNNs. Section D.1 provides the MTL experiments with single-head and multi-head architectures but with a frozen backbone. Section D.2 offers the results for the experiments with unfrozen backbones. In Section D.3, we dig deeper into investigating the optimal weights learned by each MLP layer, given that the layers are unshared across the tasks, and we measure the task similarity at each layer using the average inner product of the optimal task-specific weights. In Section D.4, we investigate the effect of the MLP depth and provide the corresponding train and test curves. Finally, in Section D.5 we offer extra experiments on CL models and demonstrate the effect of label noise on CL models.
D.1 Complementary Multi-Task Learning Experiments
Figures 7 to 13 demonstrate the full training and testing curves for training the model in an MTL setup with multi-head and single-head architectures with different backbones and datasets. As discussed before in the main text, the training error is zero for both cases in large enough models. However, the multi-task learner performs worse in the single-head architecture due to the conflict at the classification layer.
Figure 7: Comparing the single-task and multi-task performance on CIFAR-100 with pretrained ResNet-18 backbone. The top row shows the experiment with the single-head architecture and the bottom row corresponds to the multi-head architecture. The horizontal axis in all plots is the width of the MLP layer which is placed on top of the feature extractor. $\sigma$ represents the noisy portion of the training set that was corrupted by randomly switching its labels.
@@ D.1 Complementary Multi-Task Learning Experiments
D.2 Multi-Task Learning Experiments with Unfrozen Backbone
@@
Figure 14: Comparing the single-task and multi-task performance on CIFAR-100 with unfrozen ResNet-18 backbone with random initialization. The top row shows the experiment with the single-head architecture, and the bottom row corresponds to the multi-head architecture. The horizontal axis in all plots is the scale factor that controls the number of filters in convolutional layers. $\sigma$ represents the noisy portion of the training set that was corrupted by randomly switching its labels.
D.3 Investigation of the Task Optimal Weights
@@
Figure 15: Investigating the effect of weight sharing in MTL by unsharing different layers of the MLP layer and training task-specific layers. The horizontal axis is the width of the MLP layer.
Figure 16: Comparing the average inner product score (Equation 24) across different layers after unsharing the weights across the tasks. The horizontal axis is the width of the MLP layer.
D.4 Investigation of the Depth
In this section, we investigate the effect of the MLP depth on top of the pretrained backbones. Figure 17 compares different models with varying depths of the MLP. Notice that our theories are based on the assumption of a linear model, and therefore, they do not directly predict the behavior as a function of the depth. However, as discussed in Section B, the DNN performs similarly to a linear model as long as certain conditions hold. Figure 17 confirms this claim by showing similar behavior for all different models. Another interesting observation in this plot is the impact of depth over the height of the test error peak, which is a notable direction for future studies.
Figure 17: Visualizing the effect of the depth on the performance curves of the multi-head multi-task learner. The horizontal axis is the width of the MLP layers. Notice that a simple linear layer on top of the feature extractor can not fully overfit the training samples and has no tunable hyperparameter for increasing the complexity.
D.5 Complementary Continual Learning Experiments
@@
Figure 18: Visualizing the CL performance on CIFAR-100 with different values of memory budget denoted as $m$. The horizontal axis is the width of the MLP layer. The top row corresponds to the experiments with a single-head classifier and the bottom row shows the multi-head architecture.
Figure 19: Visualizing the CL performance on CIFAR-100 using an unfrozen ResNet-18 backbone and a multi-head classification head. The plot shows the performance for different values of memory budget denoted as $m$. The horizontal axis is scale factor that controls the number of filters in the ResNet-18 backbone.
@@
Figure 20: Visualizing the CL performance of multi-head models using memory size $m=50$ on CIFAR-100 with different levels of noise denoted as $\sigma$. The horizontal axis is the width of the MLP layer.
Appendix E Limitations and Future Work
@@
The next potential limitation is that our analysis is focused on linear models. Although we connected the linear models and DNNs in Appendix B, but the connection holds under certain assumptions which may be violated in different applications. Besides, we emphasize that the goal of this work is not to make formal claims regarding the performance of multi-task DNNs, rather, we tried to qualitatively connect our understanding of linear models to the observations derived experimentally from DNNs. We believe that knowing the trade-offs between different system specifications, such as the model size and number of training samples, can provide an important guideline when deploying practical models.
@@
Appendix F Infrastructures and Data Availability
We mainly used PyTorch (Paszke et al., 2019) for our implementations and datasets are freely accessible (Imagenet-R (Hendrycks et al., 2021), CUB(Wah et al., 2011), CIFAR-100 (Krizhevsky et al., 2009), and MNIST (Deng, 2012)). All experiments in the main text are reproducible on a single NVIDIA 2080 TI GeForce GPU.